# Scheduling problem of aircrafts on a same runway and dual runways

### v3

@@ VIII Simulation
In this section, we evaluate the efficiency of the proposed algorithm by comparing its computation time and objective function performance against a standard MIP solver for two scheduling problems: takeoff and landing operations on a single runway and on dual runways. For the single-runway case, we consider four scenarios with $|A|=30,40,50,60$ aircraft. For the dual-runway case, we consider scenarios with $|A|=70,80,90,100$ aircraft. The aircraft are categorized into $6$ classes based on the RECAT-EU framework $(A,B,C,D,E,F)$, with class proportions of $10\%$, $20\%$, $25\%$, $15\%$, $20\%$, and $10\%$, respectively. The minimum separation times for aircraft pairs with the same operation task are detailed in Tables IX and X. For both scheduling problems, the separation time between a takeoff aircraft and a trailing landing aircraft is set to $D_{T}=D_{P}=60$ seconds, while the separation time between a landing aircraft and a trailing takeoff aircraft is set to $T_{D}=75$ seconds for the single-runway case and $P_{D}=0$ seconds for the dual-runway case. All simulations are conducted on a computer with an AMD Ryzen 7 7840H processor (3.8 GHz, 16 GB RAM). The MIP formulations are solved using CPLEX 12.10, with a computation time limit of $600$ seconds per instance. The proposed algorithm is implemented in MATLAB R2021b.
First, we consider the aircraft scheduling problem on a single runway. The earliest landing or takeoff times for each aircraft are generated randomly. These times follow a uniform distribution within the interval $[0,T_{E}]$ minutes. The time window lengths are set to $T_{W}=30$, $45$ and $60$ minutes to evaluate the algorithms’s performance under different levels of scheduling flexibility. We evaluate three types of operations: takeoff only, landing only, and mixed operations (both takeoff and landing). The values of $T_{E}$ are set to $30$ minutes for the takeoff-only and landing-only cases and $20$ minutes for the mixed operation case. Table XI presents a comparison of the objective function values and computation times between the proposed algorithm and the MIP solver, along with the percentage gap in objective values between the two methods. The values of the objective function are denoted as NaN (Not a Number) when the MIP solver fails to find a solution within the $10$-minute time limit.
Next, we consider the aircraft scheduling problem on dual runways. The earliest landing or takeoff times for each aircraft are also generated randomly, following a uniform distribution within the interval $[0,T_{E}]$ minutes. The time windows are again set to $T_{W}=30$, $45$ and $60$ minutes. Table XII presents a comparison of the objective function values and computation times between the proposed algorithm and the MIP solver, along with the percentage gap in objective values between the two approaches.
@@ A B C D E F
60 50 Takeoff NaN 4461 600 0.83 /
50 Landing NaN 4586 600 1.11 /
40 Mixed NaN 3737 600 3.91 /
@@
From Tables XI and XII, it is evident that the proposed algorithm consistently obtains solutions within $3$ seconds, whereas the MIP solver requires the full $10$-minute time limit. Moreover, our algorithm consistently yields better objective function values than those produced by the MIP solver, with the performance gap widening as the number of aircraft increases. In addition, the runtime of the MIP solver has been extended to over one hour but no significant improvements were found in the objective function values compared to the results obtained within the $10$-minute time limit.
@@ IX Conclusions
In this paper, scheduling problems of landing and takeoff aircrafts on a same runway and on dual runways were addressed. A new theoretical framework for scheduling problem of aircrafts was established, which is completely different from the framework of mixed-integer optimization problem. Two real-time optimal algorithms are proposed for the four scheduling problems by fully exploiting the combinations of different classes of aircrafts, which can even be applied to the RECAT-EU systems. Numerical examples are presented to show the effectiveness of the algorithms. In particular, when $100$ aircrafts is considered, by using the algorithm in this paper, the optimal solution can be obtained in less than $5$ seconds, while by using the CPLEX software to solve the mix-integer optimization model, the optimal solution cannot be obtained within $1$ hour.

### v5

@@ VII-A Some necessary results
VII-B Algorithm 1
Step 2. Search for the optimal sequence of aircraft $Tcf_{11}$ and $Tcf_{12}$.
When no confusion arises, it is assumed by default that the mentioned sequence or subsequence contains the information of landing/takeoff times in Algorithms 1 and 2.
Step $i$, $i=4,5,\cdots,n$. Suppose that $\phi_{i-1}=\langle Tcf_{1},Tcf_{2},\cdots,Tcf_{i-1}\rangle$ is an optimal sequence for aircraft $Tcf_{11},Tcf_{12},\cdots,Tcf_{1(i-1)}$. Search for the optimal sequences of aircraft $Tcf_{11},Tcf_{12},\cdots,Tcf_{1i}$ based on the obtained theoretical results. The main idea is to exclude most of the non-optimal solutions and narrow down the search for the optimal solutions within a small range. The main steps are as follows.
$(i-1)$. Based on the obtained optimal sequence $\phi_{i-1}$, by a series of equivalent transformations, record all possible optimal sequences based on the bases of all the breakpoint-drift equivalent transformation sets and the takeoff-landing/landing-takeoff transitions, and represent their formed set as $\Lambda_{i-1}$ for aircraft $Tcf_{11},Tcf_{12},\cdots,Tcf_{1(i-1)}$. To find the optimal sequences of aircraft $Tcf_{11},Tcf_{12},\cdots,Tcf_{1i}$, we can first insert aircraft $Tcf_{1i}$ between any two adjacent aircraft for each sequence in the set $\Lambda_{i-1}$, denoted by $\phi_{(i-1)k}$, under the time window constraints without changing the orders of the aircraft in $\phi_{(i-1)k}$. Let $f^{i}_{inc}$ be the smallest value of the objective functions $F(\cdot)$ among all the generated new sequences.
For the landing/takeoff scheduling problem, from Theorems 2, 5 and 9, the value of the objective function $F(\cdot)$ for the landing/takeoff scheduling problem is heavily related to the separation times between breakpoint aircraft and their trailing aircraft. To obtain or record all possible optimal sequences based on the bases of all the breakpoint-drift equivalent transformation sets, emphasis should be laid on the breakpoint aircraft. For the scheduling problem of landing and takeoff aircraft on a same runway, the number of breakpoint aircraft is usually small but the number of the takeoff-landing/landing-takeoff transitions might be large, and emphasis should be laid on the takeoff-landing/landing-takeoff transitions.
The goal of this step is actually to find all possible combinations of three consecutive aircraft of different classes in all sequences of the set $\Lambda_{i-1}$ to obtain the smallest value of the objective functions $F(\cdot)$, where each combination corresponds to a time increment when the aircraft $Tcf_{1i}$ is inserted. Starting from this goal, we can reduce some computation amount by giving up some unnecessary operations. For example, since we are concerned about only the time increment when the aircraft $Tcf_{1i}$ is inserted into $\phi_{(i-1)k}$, we need not to repeatedly consider the insertion of aircraft $Tcf_{1i}$ into the same class combination of aircraft and the same separation times, and only focus on the different class combination of aircraft or different separation times, which might significantly reduce the computation amount.
Time increment would be used frequently in our algorithm. To illustrate its role specifically, we give an example. Generate two landing sequences $\phi_{1}=\langle Tcf_{1},Tcf_{4},Tcf_{2},Tcf_{3}\rangle$ and $\phi_{2}=\langle Tcf_{1},Tcf_{2},Tcf_{4},Tcf_{3}\rangle$ by inserting aircraft $Tcf_{4}$ into a sequence $\phi_{0}=\langle Tcf_{1},Tcf_{2},Tcf_{3}\rangle$. The time increment of the separation time between aircraft $Tcf_{1}$ and $Tcf_{2}$ in $\phi_{1}$ with respect to $\phi_{0}$ is $\Delta_{1}=S_{14}(\phi_{1})+S_{42}(\phi_{1})-S_{12}(\phi_{0})$ and the time increment of the separation time between aircraft $Tcf_{1}$ and $Tcf_{2}$ in $\phi_{1}$ with respect to $\phi_{0}$ is $\Delta_{2}=S_{24}(\phi_{2})+S_{43}(\phi_{2})-S_{23}(\phi_{0})$. If $S_{23}(\phi_{1})=S_{23}(\phi_{0})$ and $S_{12}(\phi_{2})=S_{12}(\phi_{0})$, then $F(\phi_{2})-F(\phi_{1})=\Delta_{2}-\Delta_{1}$.
[…]
$(i-3).$ According to the elements of the set $F_{inc}$, adjust the aircraft orders in the sequence $\phi_{(i-1)k}$ and insert the aircraft $Tcf_{1i}$ to generate a new sequence $\bar{\phi}_{ik}$ under the condition $F(\bar{\phi}_{ik})<f^{i}_{inc}$ such that $\bar{\phi}_{ik}$ contains the subsequence $\langle Tcf_{p_{1}},Tcf_{k_{1}},Tcf_{1i},Tcf_{j_{1}}\rangle$ and $\langle p_{1},{k_{1}},{j_{1}}\rangle\in F_{inc}$. Minimize $F(\bar{\phi}_{ik})$ with the subsequence $\langle Tcf_{p_{1}},Tcf_{k_{1}},Tcf_{1i},Tcf_{j_{1}}\rangle$ unchanged based on the obtained theoretical results. Compare the values of all possible $F(\bar{\phi}_{ik})$ so as to find the optimal sequence of aircraft $Tcf_{11},Tcf_{12},\cdots,Tcf_{1i}$.
To illustrate this step in a more specific way, consider a group of landing aircraft with $T_{0}=60$ and $\delta=8$ under the minimum separation time standards at Heathrow Airport. From the previous analysis in (i-2), the inequality (20) is equivalent to
$Y_{k(1i)}+Y_{(1i)j}-Y_{kj}<f^{i}_{inc}-F(\phi_{i-1}).$ (21)
If $f^{i}_{inc}-F(\phi_{i-1})=T_{0}$, by simple calculations, the inequality (21) holds when $(cl_{k},cl_{1i},cl_{j})=(2,5,4),(3,6,5),(4,6,5)$. Clearly, when $(cl_{k},cl_{1i},cl_{j})=(2,5,4),(3,6,5),(4,6,5)$, each aircraft $Tcf_{k}$ is a breakpoint aircraft and the aircraft $Tcf_{1i}$ is needed to be inserted to be between aircraft $Tcf_{k}$ and $Tcf_{j}$. To generate a subsequence $\langle Tcf_{k},Tcf_{1i},Tcf_{j}\rangle$, we can focus on the breakpoint aircraft and their trailing aircraft, and based on the theoretical results in previous sections, by adjusting the aircraft orders properly, the optimal sequence can be found under the condition that $F(\bar{\phi}_{ik})<f^{i}_{inc}$ in an easy way. If $f^{i}_{inc}-F(\phi_{i-1})=T_{0}+\delta$, the generation of a subsequence $\langle Tcf_{k_{1}},Tcf_{1i},Tcf_{j_{1}}\rangle$ with $\langle{k_{1}},{j_{1}}\rangle\in F_{inc}$ would result in a new breakpoint aircraft or changes of the separation times between existing breakpoint aircraft and their trailing aircraft which might increase the value of the objective function $F(\cdot)$. Note for the separation times between breakpoint aircraft and their trailing aircraft that $T_{45}=T_{0}+\delta$ and $T_{ij}\geq 1.5T_{0}$ for all other $1\leq i<j\leq 6$. We can generate the whole sequence based on the subsequence $\langle Tcf_{k_{1}},Tcf_{1i},Tcf_{j_{1}}\rangle$ according to time increments of the new breakpoint aircraft or changes of the separation times between breakpoint aircraft and their trailing aircraft under the condition that $F(\bar{\phi}_{ik})<f^{i}_{inc}$, and the optimal sequence can also be found in an easy way. Besides these two cases, the case of $f^{i}_{inc}-F(\phi_{i-1})>T_{0}+\delta$ as well as the takeoff sequences and mixed landing and takeoff sequences on a same runway can be discussed in a similar easy way.
From the above illustrations, we can first move the aircraft $Tcf_{p_{1}},Tcf_{k_{1}},Tcf_{j_{1}}$ to proper places by equivalent transformations including breakpoint-drift equivalent transformations and then use the extraction operation and insertion operation to generate the sequence $\bar{\phi}_{ik}$ with a focus on the breakpoint aircraft, their trailing aircraft and the landing-takeoff/takeoff-landing transitions as well as the properties of class-monotonically-decreasing sequences on the basis of the optimal sequence.
Since $Y_{p_{1}k_{1}}+Y_{k_{1}(1i)}+Y_{(1i)j_{1}}-Y_{p_{1}j_{1}}<f^{i}_{inc}-F(\phi_{i-1})$, there is no utilizable subsequence $\langle Tcf_{p_{1}},Tcf_{k_{1}},Tcf_{j_{1}}\rangle$ in each $\phi_{(i-1)k}$ for $Tcf_{1i}$ to be inserted between aircraft $Tcf_{k_{1}}$ and $Tcf_{j_{1}}$. So, the generation of a subsequence $\langle Tcf_{p_{1}},Tcf_{k_{1}},Tcf_{j_{1}}\rangle$ would increase the value of the objective function $F(\cdot)$ for aircraft $Tcf_{11},Tcf_{12},\cdots,Tcf_{1(i-1)}$. Moreover, the optimal insertion of $Tcf_{1(i-1)}$ might usually generate a new breakpoint/landing-takeoff/takeoff-landing transition or change an existing separation time between a breakpoint aircraft and its trailing aircraft with some necessary aircraft order adjustments. These properties and related discussions in previous sections can be used to reduce the computation amount of Algorithm 1.
Remark 38
Since different classes might correspond to different time increments, we can search for the desired aircraft of proper classes in $F_{inc}$ in ascending order of time increments to minimize the value of the objective function $F(\cdot)$. Moreover, note that the separation times between aircraft of different classes are time-invariant and the number of the total aircraft classes is not very large. The computation cost of time increments is also not very large and can even be calculated out offline.
Remark 39
It should be noted that in Algorithm 1, when there is no resident-point aircraft, the optimal sequence $\phi_{i}$ satisfies that $F(\phi_{i-1},Sr(\phi_{i-1}))\leq F(\phi_{i},Sr(\phi_{i}))\leq F(\phi_{i-1},Sr(\phi_{i-1}))+Y_{(i-1)i}$ for the landing/takeoff scheduling problem, and the landing and takeoff scheduling problem on a same runway, where $T_{0}\leq Y_{(i-1)i}\leq 3T_{0}$. For example, consider a sequence $\phi_{0}=\langle Tcf_{1},Tcf_{2}\rangle$, where $Tcf_{1},Tcf_{2}$ are both landing aircraft and $Y_{12}>T_{D}+D_{T}$. Generate a new sequence $\phi_{1}$ by inserting a takeoff aircraft $Tcf_{3}$ to be between $Tcf_{1}$ and $Tcf_{2}$, where $S_{13}(\phi_{1})=T_{D}$ and $S_{12}(\phi_{1})=S_{12}(\phi_{0})$. It is clear that $F(\phi_{1})-F(\phi_{0})=0$.
Remark 40
Remark 41
VII-C Algorithm 2
Algorithm 2. Suppose that there are totally $n+m$ aircraft composed of $n$ landing aircraft and $m$ takeoff aircraft, denoted by $Tcf_{1},Tcf_{2},\cdots,Tcf_{n+m}$, and there is at least a feasible sequence for them to land/takeoff without conflicts. Let $Tcf^{0}_{1},Tcf^{0}_{2},\cdots,Tcf^{0}_{n}$ denote all the landing aircraft, and $Tcf^{1}_{1},Tcf^{1}_{2},\cdots,Tcf^{1}_{m}$ denote all the takeoff aircraft. The main idea of this algorithm is also to exclude most of the non-optimal solutions and narrow down the search for the optimal solutions within a small range. The main steps are as follows.
The goal of recording all possible optimal sequences is to make preparations to generate all possible block subsequences in Steps 3 and 4. Therefore, we can reduce some computation amount according to the needs of Steps 3 and 4.
When our algorithm is applied, in the landing/takeoff sequence of each block subsequence, we consider only the case of one breakpoint aircraft and the case of two or more breakpoint aircraft in a block subsequence can be dealt with by combing the operations of extraction and insertion to decrease the value of the objective functions based on the analysis of the case of one breakpoint aircraft for block subsequences. Or alternatively, the case of two or more breakpoint aircraft in a block subsequence can be dealt with offline by using the same method for the case of one breakpoint aircraft.
As discussed in Sec. VI, if the number of breakpoint aircraft in a block subsequence is fixed, the number of all possible block subsequences of different class combinations is not large according to the minimum separation time standards at Heathrow Airport and for the RECAT-EU system, when resident-point aircraft are not taken into account.
(2.2) Focus on combinations of different block subsequences, and study the landing/takeoff time increments of any two given consecutive block subsequences, where the last two aircraft of the first block subsequence is the first two aircraft of the second block subsequence. According to the landing/takeoff time increments, classify all possible two consecutive subsequences into several sets.
In our simulations for the case of $F(\Phi_{a})\geq F(\Phi_{b})$, the combination of a D-block subsequence and a T-block subsequence is commonly used, which has the form of $\langle DB,TB\rangle$ where $DB$ denotes a D-block subsequence and $TB$ denotes a T-block subsequence, the first two aircraft of $DB$ have the same landing/takeoff time, and the last two aircraft of $TB$ have the same landing/takeoff time (see Tables III and I for related discussions). The landing time increment of D-block subsequence in such a combination is a key quantity that needs to be studied for the optimal sequence.
Under Assumption 17, the sequence $\Phi_{c}$ can be divided into a series of block subsequences. Since $F(\Phi^{c}_{LA},Sr_{LA}(\Phi_{c}))=F(\Phi_{a})$, we can use a dynamic programming approach to find the optimal sequence $\Phi_{c}$ based on the sets defined in Step 2. It should be noted that at each step of the dynamic programming approach, we need consider all the possible block subsequences that can decrease the objective function $F(\cdot)$ of the whole sequence, and if the generation of one block subsequence might disrupt the existing block subsequences, we can first generate the block subsequence and then use it as a basis to generate other block subsequences in an optimal way.
Here, it should be noted that each sequence of $\Lambda_{a}$ is required to be considered. Actually, we usually need to consider only the impact of the local differences of the sequence from other sequences instead of the whole sequence, which might not need a lot of the computation amount.
Step 4. (4.1) The main idea of this step is to generate a solution $(\Phi_{c4},Sr(\Phi_{c4}))$ such that $\Phi^{c4}_{LA}\in\Lambda_{a}$, and search for the optimal solutions for the optimization problem (6) under the condition that $F(\Phi_{a})\leq F(\Phi^{c4}_{LA})\leq F(\Phi_{c3})$. Note that the set $\Lambda_{a}$ used here does not contain the information of the landing/takeoff times.
Since different D-block subsequences might have different landing time increments, we can search for the optimal solutions for the optimization problem (6) according to the sum of the landing time increments of D-block subsequences of the sequence $\Phi_{c4}$.
If the sum of the landing time increments of D-block subsequences of $\Phi_{c4}$ is larger than $F(\Phi_{c3})-F(\Phi_{a})$, then it follows that $F(\Phi_{c4})>F(\Phi_{c3})$, implying that $(\Phi_{c4},Sr(\Phi_{c4}))$ is not an optimal solution for the optimization problem (6).
When D-block subsequences of $\Phi_{c4}$ are considered, some necessary aircraft order adjustments might be needed to be imposed on the sequence $\Phi_{c4}$ to fully consider all the possibilities of the block subsequences to reduce the value of the objective function $F(\cdot)$. For example, suppose that $\phi_{LA}^{0}=\langle Tcf_{1},Tcf_{2},Tcf_{3},Tcf_{4},Tcf_{5}\rangle$ is a landing subsequence of a sequence $\phi_{0}$, and $\phi_{0}$ is composed of a D-Block subsequence $DB_{1}$ and a T-Block subsequence $TB_{1}$ with $Tcf_{1},Tcf_{2}\in DB_{1}$ and $Tcf_{2},Tcf_{3},Tcf_{4},Tcf_{5}\in TB_{1}$. According to the definition of D-block subsequence, $S_{12}>Y_{12}$. If $Y_{15}>Y_{12}$ and the objective function value of the landing sequence can be decreased by exchanging the orders of $Tcf_{2}$ and $Tcf_{5}$, the D-block subsequences of the sequence $\phi=\langle Tcf_{1},Tcf_{5},Tcf_{3},Tcf_{4},Tcf_{2}\rangle$ are needed to be considered.
Different from the approach described above, there is another search approach to generate the optimal sequence directly according to block subsequences and the combinations of block subsequences discussed in Step 2, which might have more computation amount.
Remark 42
The main idea of Algorithm 2 is to decompose the aircraft sequence into several block subsequences and fully explore combinations of the block subsequences along an optimal landing/takeoff subsequence to consider the optimization problem (6). Note that the number of all possible block subsequences of different class combinations is not large according to the minimum separation time standards at Heathrow Airport and for the RECAT-EU system, when resident-point aircraft are not taken into account. The use of block subsequences might significantly reduce the computation amount of the algorithm.
Remark 43
The complexities of Algorithms 1 and 2 are heavily related to the aircraft number, the aircraft classes, and the constraints of the aircraft time windows. But the algorithms essentially exhibit the features of polynomial algorithms and can be applied in actual systems in real time, and based on these two algorithms, the optimal solutions for the optimization problem (6) can be obtained.
@@ VIII Simulation
In this section, we evaluate the efficiency of the proposed algorithms by comparing its computation time and objective function performance against a standard MIP solver for two scheduling problems: takeoff and landing operations on a single runway and on dual runways. For the single-runway case, we consider four scenarios with $|A|=30,40,50,60$ aircraft. For the dual-runway case, we consider scenarios with $|A|=70,80,90,100$ aircraft. The aircraft are categorized into $6$ classes based on the RECAT-EU framework $(A,B,C,D,E,F)$, with class proportions of $10\%$, $20\%$, $25\%$, $15\%$, $20\%$, and $10\%$, respectively. The minimum separation times for aircraft pairs with the same operation task are detailed in Tables IX and X. For both scheduling problems, the minimum separation times between a takeoff aircraft and a trailing landing aircraft are set to $D_{T}=D_{P}=60$ seconds, while the minimum separation time between a landing aircraft and a trailing takeoff aircraft is set to $T_{D}=75$ seconds for the single-runway case and $P_{D}=0$ seconds for the dual-runway case. All simulations are conducted on a computer with an AMD Ryzen 7 7840H processor (3.8 GHz, 16 GB RAM). The MIP formulations are solved using CPLEX 12.10, with a computation time limit of $600$ seconds per instance. The proposed algorithms are implemented in MATLAB R2021b.
First, we consider the aircraft scheduling problem on a single runway. The earliest landing or takeoff times for each aircraft are generated randomly. These times follow a uniform distribution within the interval $[0,T_{E}]$ minutes. The time window lengths are set to $T_{W}=30$, $45$ and $60$ minutes to evaluate the algorithms’s performance under different levels of scheduling flexibility. We evaluate three types of operations: takeoff only, landing only, and mixed operations (both takeoff and landing). The values of $T_{E}$ are set to $30$ minutes for the takeoff-only and landing-only cases and $20$ minutes for the mixed operation case. Table XI presents a comparison of the objective function values and computation times between Algorithm 1 and the MIP solver, along with the percentage gap in objective values between the two methods. The values of the objective function are denoted as NFS when the MIP solver fails to find a solution within the $10$-minute time limit.
Next, we consider the aircraft scheduling problem on dual runways. The earliest landing or takeoff times for each aircraft are also generated randomly, following a uniform distribution within the interval $[0,T_{E}]$ minutes. The time windows are again set to $T_{W}=30$, $45$ and $60$ minutes. Table XII presents a comparison of the objective function values and computation times between Algorithm 2 and the MIP solver, along with the percentage gap in objective values between the two approaches.
@@ A B C D E F
60 50 Takeoff NFS 4461 600 0.83 /
50 Landing NFS 4586 600 1.11 /
40 Mixed NFS 3737 600 3.91 /
@@
From Tables XI and XII, it is evident that the proposed algorithms consistently obtains solutions within $5$ seconds, whereas the MIP solver requires the full $10$-minute time limit. Moreover, our algorithmx consistently yield better objective function values than those produced by the MIP solver, with the performance gap widening as the number of aircraft increases. In addition, the runtime of the MIP solver has been extended to over one hour but no significant improvements were found in the objective function values compared to the results obtained within the $10$-minute time limit.
@@ IX Conclusions
In this paper, scheduling problems of landing and takeoff aircraft on a same runway and on dual runways were addressed. A new theoretical framework for scheduling problem of aircraft was established, which is completely different from the framework of mixed-integer optimization problem. Two real-time optimal algorithms are proposed for the four scheduling problems by fully exploiting the combinations of different classes of aircraft, which can even be applied to the RECAT-EU systems. Numerical examples are presented to show the effectiveness of the algorithms. In particular, when $100$ aircraft are considered, by using the algorithm in this paper, the optimal solution can be obtained in less than $5$ seconds, while by using the CPLEX software to solve the mix-integer optimization model, the optimal solution cannot be obtained within $1$ hour.
@@ X Reference
[30] M., Charles, J. Peters, and P. Chorora, “Validation of the time based separation concept at London Heathrow Airport,” Tenth USA/Europe Air Traffic Management Research and Development Seminar, 2013.